OpenStudy (accidentalaichan):

What is the exact value of the expression 7 sqrt 8 + 2 sqrt 12 - 3 sqrt 2?

OpenStudy (luigi0210):

\(\Large 7 \sqrt{8}+2\sqrt{12}-3\sqrt{2}=7\sqrt{4(2)}+2\sqrt{6(2)}-3\sqrt{2}\) \(\Large (7*2)\sqrt{2}+4\sqrt{3}-3\sqrt{2}\) \(\Large (14\sqrt{2}-3\sqrt{2})+4\sqrt{3}\)

OpenStudy (accidentalaichan):

Thanks, Jesus! :3